$SPAD/src/input defintef.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool defintef.output
)set message test on
)set message auto off
)clear all

———-

Evans and Johnson, in their “Uses of technology in the mathematics curriculum” workbook list a number of “integrals that your computer can’t do”.

int(sin(x)^3/(sin(x)^3+cos(x)^3),x=0..Pi/2)

is one of them.

— * —

--S 1 of 8
sin(x)**3/(sin(x)**3+cos(x)**3)
--R
--R
--R                   3
--R             sin(x)
--R   (1)  -----------------
--R              3         3
--R        sin(x)  + cos(x)
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 8
integrate(%, x = 0..%pi/2, "noPole")
--R
--R
--R        2log(16) - 4log(4) + 3%pi
--R   (2)  -------------------------
--R                    12
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 2
--S 3 of 8
x**2/(1+x**3)

--R
--R
--R           2
--R          x
--R   (3)  ------
--R         3
--R        x  + 1
--R                                          Type: Fraction(Polynomial(Integer))
--E 3
--S 4 of 8
integrate(%, x=0..%plusInfinity)
--R
--R
--R   (4)   + infinity
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 4
--S 5 of 8
exp(-x**2)*log(x)**2
--R
--R
--R             2
--R          - x       2
--R   (5)  %e    log(x)
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 8
integrate(%, x=0..%plusInfinity)
--R
--R
--R         _1             1     _1         1 2
--R        | (-)polygamma(1,-) + | (-)digamma(-)
--R           2             2       2         2
--R   (6)  --------------------------------------
--R                           8
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 6
--S 7 of 8
x * asin(x/(x+1))
--R
--R
--R                 x
--R   (7)  x asin(-----)
--R               x + 1
--R                                                    Type: Expression(Integer)
--E 7

--S 8 of 8
integrate(%, x=0..1)
--R
--R
--R          +-+
--R        3\|3  - 4
--R   (8)  ---------
--R            6
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 8
)spool
)lisp (bye)

———-

References^{[1] nothing}

6